Solve for x and y
2x + 2y=8
6x - y=10

Answer by mananth(16949)

2 x + 2 y = 8 .............1
6 x -1 y = 10 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 2
2 x 2 y = 8
12 x -2 y = 20
Add the two equations
14 x = 28
/ 14
x = 2
plug value of x in (1)
2 x + 2 y = 8
4 + 2 y = 8
2 y = 8 -4
2 y = 4
y = 2